Determine the horizontal and vertical components of reaction on the member at the pin A, and the normal reaction at the roller B.


1
Expert's answer
2021-09-06T00:53:23-0400

Take moment about A.

"\Sigma M =O"

NBcos30(6)-NBsin30(2)-750(3)=0

5.19NB-NB-2250=0

NB=536.22 lb


equilibrium of forces along x-direction

"W\Sigma Fx=0"

Ax=NBsin30

Ax=536.22sin30

Ax=268.11 lb


equilibrium of forces along y-direction

"W\Sigma Fy=0"

Ay+NBCos(30)-750=0

Ay+536.22Cos(30)-750=0

Ay=285.61 lb